OpenStudy (anonymous):

Object A with a mass of 1.5 kilograms is moving with a velocity of +11.2 meters/second (moving in the +x direction). It has a perfectly elastic collision with stationary object B that has a mass of 4.5 kilograms. After the collision, object B travels with a velocity of +5.6 meters/second. What is the final velocity for object A?

OpenStudy (anonymous):

use conservation of momentum

OpenStudy (anonymous):

What he said means that the amount of Kinetic Energy will be the same after impact, just distributed differently. this helped me: Object A's kinetic energy transfers into object B's kinetic energy. A's energy - B's energy = kinetic energy of object A after the collision. Kinetic energy can be calculated using the following equation, where m=mass (kg); v=velocity (m/s); and E=energy (J) E=1/2*m*v^2 E(A) - E(B) = E(A after the collision) 1/2*1.5kg*(11.2m/s)^2 - 1/2*4.5kg*(5.6m/s)^2 = 23.52 J (J=joule) There's still some energy left, so object A didn't change direction. Object A's kinetic energy is now 23.52 J. We can now calculate it's velocity. 1/2*1.5kg*x^2=23.52 J | /(1/2*1.5) x^2=23.52/(1/2*1.5) x^2=31.36 | square root x=+-sqrt(31.36) x=+-5.6 (m/s) we choose +5.6 m/s only, as object A didn't change its direction. Answer: C: +5.6m/s in the same direction.

OpenStudy (anonymous):

KE=mv^2 So you add the two to get TOTAL KE